OpenStudy (anonymous):

Subtract. –0.61 – (–7.74) A. –8.35 B. –7.13 C. 7.13 D. 8.35

OpenStudy (ahvei):

C

OpenStudy (anonymous):

ty

OpenStudy (ahvei):

yw

OpenStudy (texaschic101):

-0.61 - (-7.74) = -0.61 + 7.74 = 7.13